OpenStudy (anonymous):

h=4vt-19t^2 solve for v

OpenStudy (anonymous):

I am getting h/4t + 19t/4.

OpenStudy (anonymous):

thanks, can you explaing

OpenStudy (solomonzelman):

You have to isolate the v! h = 4vt-19t^2 -4vt = -h -19t^2 -4vt / (-4t) = ( -h -19t^2 ) / (-4t)

OpenStudy (solomonzelman):

v =( -h -19t^2 ) / (-4t)

OpenStudy (solomonzelman):

OR h/ ( 4t ) + 19t^2 / 4t Understand?... Need explanations?

OpenStudy (solomonzelman):

x3pufflix3 was correct!